# DAG-Plan: Generating Directed Acyclic Dependency Graphs for Dual-Arm Cooperative Planning

### v2

@@ A.3 Reinforcement Learning
$obj\_grasped$ Grasp target object $1\enspace if\enspace\text{grasped}(target)\enspace else\enspace 0$
$obj\_in\_obj$ Put object into target object $1\enspace if\enspace\text{in}(obj,target)\enspace else\enspace 0$
$obj\_on\_obj$ Put object onto target object $1\enspace if\enspace\text{on}(obj,target)\enspace else\enspace 0$
$joint\_qpos$ Move joint to target joint position $\frac{\text{clip}(pos_{cur},pos_{min},pos_{max})-pos_{min}}{pos_{max}-pos_{min}}$
$ee\_pos$ Move end-effector to target object $\frac{1}{2}(1-\frac{\text{dis}(hand,target)}{init\_dis}+1-\text{tanh}(\text{dis}(hand,target)))$
$ee\_height$ Keep both end-effectors at same height $1-\frac{\text{min}((\text{abs}(z_{right}-z_{left})),0.05)}{0.05}$
$obj\_cut$ Cut the target object with object (knife) $\frac{1}{2}(obj\_grasped(obj)+obj\_on\_obj(obj,target))$
$obj\_pour$ Pour liquid in object into target object $\frac{1}{3}(obj\_grasped(obj)+obj\_horizon(obj)+obj\_above(obj,target))$
Table 6: Reward and success condition of Sub-tasks. Success condition is marked in bold.
Sub-task Pseudocode
$grasp<target\_obj>$ $1.0\times ee\_pos(target)+1.0\times\mathbf{obj\_grasped(target)}$
$pour<obj>into<target\_obj>$ $1.0\times\mathbf{obj\_pour(obj,target)}$
$cut<target\_obj>with<obj>$ $1.0\times\mathbf{obj\_cut(obj,target)}$
$put<obj>into<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_in(obj,target)}$
$put<pot>into<sink>$ $10.0\times ee\_both(sink)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_in(pot,sink)}+1.0\times\mathbf{obj\_horizon(pot)}$
$put<obj>onto<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_on(obj,target)}$
$put<pot>onto<stove\_body>$ $10.0\times ee\_both(stove\_body)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_on(pot,stove\_body)}+1.0\times\mathbf{obj\_horizon(pot)}$
$open<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$close<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$switch\enspace on<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$switch\enspace off<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
Appendix B Dual-arm Kitchen Benchmark Details
B.1 Sub-tasks Action Capabilities
The action capabilities of sub-tasks in the Dual-arm Kitchen Benchmark are listed in Table 7.
Table 7: Sub-tasks action capabilities. Different actions correspond to the node types occupy, tool use, release, and operate.
Sub-task Description
$grasp<target\_obj>$ Grasp target object
$pour<obj>into<target\_obj>$ Pour the liquid in object into target object
$cut<target\_obj>with<obj>$ Cut target object which is food with object which is a knife
$put<obj>into<target\_obj>$ Put object in hand into target object
$put<obj>onto<target\_obj>$ Put object in hand onto target object
$open<target\_obj>$ Open target object, such as a door, a drawer floor, a faucet
$close<target\_obj>$ Close target object, such as a door, a drawer floor, a faucet
$switch\enspace on<target\_obj>$ Switch on target object which is an electric device
$switch\enspace off<target\_obj>$ Switch off target object which is an electric device
B.2 Full Sub-tasks List
The sub-tasks of each task in the Dual-arm Kitchen Benchmark are listed in Table 9.
Table 8: Sub-tasks of Dual-arm Kitchen Benchmark.
Index Instruction Sub-tasks
Task 1 Put the apple and bread onto the plate Grasp apple, Put apple onto plate, Grasp bread, Put bread onto plate
Task 2 Place the apple on the plate and toast the bread Grasp apple, Put apple onto plate, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
Task 3 Juice the apple and toast the bread Grasp apple, Put apple into juicer container, Switch on juicer, Switch off juicer, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
Task 4 Wash the cup and the bowl Open dishwasher door, Grasp cup, Put cup into dishwasher rack, Grasp bowl, Put bowl into dishwasher rack, Close dishwasher door, Switch on dishwasher, Switch off dishwasher
Task 5 Heat the soup and put the tin on the table Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp tin, Put tin onto table, Close refrigerator cooler door
Task 6 Heat the soup and pour a cup of cola Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp cola bottle, Pour cola bottle into cup, Put cola bottle into refrigerator cooler body, Close refrigerator cooler door
Task 7 Put the apple and the pear into the bowl Open cabinet left door, Grasp bowl, Put bowl onto table, Grasp apple, Put apple into bowl, Grasp pear, Put pear into bowl
Task 8 Cut the carrot and red bell pepper Open cabinet left door, Grasp knife, Grasp carrot, Put carrot onto cutting board, Cut carrot with knife, Grasp red bell pepper, Put red bell pepper onto cutting board, Cut red bell pepper, Put knife onto cutting board
Task 9 Make a pot of soup Open cabinet right door, Grasp pot, Put pot into sink, Open faucet, Close faucet, Grasp pot, Put pot onto stove body, Switch on stove, Grasp carrot, Put carrot into pot, Grasp red bell pepper, Put red bell pepper into pot, Switch off stove
B.3 Assets of Dual-arm Kitchen Benchmark
The assets of each task in the Dual-arm Kitchen Benchmark are listed in Table 9. These assets were sourced from Mobility-Partnet [^4] https://sapien.ucsd.edu/browse, BlenderKit [^5] https://www.blenderkit.com, and Sketchfab [^6] https://sketchfab.com. We have modified them and constructed the corresponding URDF models.
Table 9: Assets of Dual-arm Kitchen Benchmark.
Index Instruction Assets
Task 1 Put the apple and bread onto the plate table, apple, bread, juicer, toaster
Task 2 Place the apple on the plate and toast the bread table, apple, bread, juicer, toaster
Task 3 Juice the apple and toast the bread table, apple, bread, juicer, toaster
Task 4 Wash the cup and the bowl table, dishwasher, microwave, cup, bowl, refrigerator cooler
Task 5 Heat the soup and put the tin on the table table, dishwasher, microwave, tin, bowl, refrigerator cooler
Task 6 Heat the soup and pour a cup of cola table, dishwasher, microwave, cola bottle, bowl, refrigerator cooler
Task 7 Put the apple and the pear into the bowl table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, apple, pear
Task 8 Cut the carrot and red bell pepper table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
Task 9 Make a pot of soup table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
Appendix C Planning Evaluation Analysis
To further illustrate the differences between our approach and the baselines, we analyzed the planning evaluation of TP-S, TP-D, and DAG-Plan for tasks 5 and 8. We validated the generated plans using Planning Domain Definition Language (PDDL) written by robotics experts for the corresponding tasks. The plans generated by TP-S and TP-D were executed sequentially in the PDDL environment. For DAG-Plan, due to the lack of environmental information in this evaluation, only the first check was conducted, and the first candidate node from the candidate list was selected for execution.
C.1 Planning Evaluation of Task 5
We provide the successful plan for each method in Figure 6 and the failed plan for TP-D in Figure 7. We further analyzed the failed plans of TP-D. TP-D made an error, incorrectly using the left hand, which was holding the tin, to close the refrigerator cooler door. In addition, TP-D is missing the sub-task of closing the microwave door.
C.2 Planning Evaluation of Task 8
We provide the successful plan for each method in Figure 8 and the failed plan for each method in Figure 9. We further analyzed the failed plans of each method. TP-S did not follow the rules in this task, failing to specify which hand to use for using the knife. TP-D made an error due to the complexity of dual-arm tasks, incorrectly using the left hand, which was not holding the knife, to cut. DAG-Plan did not follow the rules initially, generating an incomplete DAG, but after reflecting to LLMs for correction, it ultimately produced the same DAG as shown in Figure 8.
C.3 Failed Plan Analysis of TP-D
We further analyzed the failed plans of TP-D on the Dual-arm Kitchen Benchmark, as shown in Figure 10. It is evident that directly generating sequence dual-arm plans using LLMs often leads to errors, such as the robot holding an object in its hand while attempting to manipulate another object. This type of sequence temporal dependency works well for generating single-arm plans but struggles with the complexity of dual-arm planning, making it difficult to produce correct dual-arm plans. In contrast, DAG-Plan imposes strict constraints with occupy-release pair can be performed once the robot holds an object. This ensures that such errors are prevented, maintaining the conciseness of the plan.
Figure 6: Successful plan for each method of task 5.
Figure 7: Failed plan for each method of task 5. Errors are highlighted in red boxes.
Figure 8: Successful plan for each method of task 8.
Figure 9: Failed plan for each method of task 8. Errors are highlighted in red boxes.
Figure 10: Failed plan for TP-D of task 2, task 4, and task 9. Errors are highlighted in red boxes.
Appendix D Execution Process Analysis
D.1 Execution Process for DAG-Plan of Task 5
We provide the full execution process for DAG-Plan of task 5 in Figure 16.
(a) Stage 1: Open microwave door (left) and Open refrigerator cooler door (right)
(a) Stage 2: Grasp tin (left) and Grasp bowl (right)
(a) Stage 3: Put tin onto table (left) and Put bowl into microwave body (right)
(a) Stage 4: Close microwave door (left) and Close refrigerator cooler door (right)
(a) Stage 5: Switch on microwave (left)
(a) Stage 6: Switch off microwave (left)
Figure 16: Execution process for DAG-Plan of task 5.
D.2 Execution Process for DAG-Plan of Task 8
We provide the full execution process for DAG-Plan of task 5 in Figure 23.
(a) Stage 1: Grasp red bell pepper (left) and Grasp carrot (right)
(a) Stage 2: Put red bell pepper onto cutting board (left) and Put carrot onto cutting board (right)
(a) Stage 3: Open cabinet left door (right)
(a) Stage 4: Grasp knife (right)
(a) Stage 5: Cut carrot with knife (right)
(a) Stage 6: Cut red bell pepper with knife (right)
(a) Stage 7: Put knife onto cutting board (right)
Figure 23: Execution process for DAG-Plan of task 8.
D.3 Failed Execution Analysis of TP-D
Since TP-D generates dual-arm plans without the ability to adjust hand movements in real-time based on environmental information, some plans that pass PDDL validation still fail in the physical simulation environment. For instance, as shown in Figure 24, "Open refrigerator cooler door (left)" and "Open microwave door (right)" cannot be executed due to the hands crossing each other’s paths.
Figure 24: Failed Execution Process of TP-D.
Appendix E Full Prompts
E.1 Our Method: Prompt for DAG-Plan
{mdframed}
[backgroundcolor=azure,linewidth=0pt] You are a helpful assistant to plan a definite task into Directed Acyclic Graph (DAG) for a dual-arm robot.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Create a DAG based on the task name, considering the task complete when all nodes are executed. The task can be complex but should be broken into simple, reasonable nodes.
2) The node name must be selected in command API and follow the description.
3) When moving an object, confirm it clearly specify its destination, e.g., ‘put "obj_name" into/onto "obj_name"’
4) Synthesize the ’edge’ of a node, a list of node’s indices the precondition of this node. When generating edges, please match the current node with all other nodes to determine if there are any dependencies.
5) Your generated edges are not only the preconditions for the node to be executed, but also preconditions to complete the task objectives. For example, in task: fill cup with coffee, node: switch on coffee machine is a executed precondition for node: switch off coffee machine to execute, and additionally, node: put cup into coffee machine is also a task precondition for node: switch off coffee machine. If this condition is missing, switch off coffee machine may executed before put cup into coffee machine, the container will not filled with coffee, and the task cannot be completed.
6) Each node has a node type. "occupy" means the hand remains occupied after the node such as node: grasp "apple", and "release" means the object is released after this node such as node: put "apple" into "drawer", and "operate" means robot manipulate joint such as node: open "drawer", and "tool use" means robot utilize obj which handling in hand such as node: cut "apple" with "knife".
7) Each occupy node must have a corresponding release node which means the obj must holding in hand before put obj.
8) When two hands are needed to grasp or put an object, use a two-handed node directly instead of two one-handed nodes. For example, use node:grasp pot instead node: grasp pot_left_handle and node: grasp pot_right_handle.
9) Ensure the number of arms used matches the "arm_num" term. The "occupy" type nodes in the edge list should not exceed two.
10) Enclose all object references in quotes "". These must be in obj_names. Do not introduce new objects or remove any.
11) You need to generate as few nodes as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any node that is not relevant to the completion of the task should not be present.
12) After the graph is generated, create a task complete node.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
  node type: operate
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
{mdframed}[backgroundcolor=azure, linewidth=0pt]   - switch_off <target-object>
  node type: operate
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
  node type: release
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour <object> into <target_object>
  node type: tool use
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
  node type: tool use
* The following command allows robot to take an <target-object>:
  command syntax:
  grasp <target-object>
  node type: occupy
Here’s an example input and response:
INPUT:
task: put apple into drawer
obj_names:  ["apple", "drawer_shell","drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: "apple" is on "table".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints: ["drawer_joint"]
        spatial_relationship: "drawer" is on "table". "drawer" is closed now.
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
        "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
{mdframed}[backgroundcolor=azure, linewidth=0pt]     "table":
        type: actor
        fixed: True
        spatial_relationship: "table" on the ground.
        description: None
RESPONSE:
nodes:
    node_1:
        name: open "drawer_floor"
        arm_num: 1
        edge: []
        type: operate
        description: Open the drawer to access inside.
    node_2:
        name: grasp "apple"
        arm_num: 1
        edge: []
        type: occupy
        description: Grasp the apple on the table.
    node_3:
        name: put "apple" into "drawer_floor"
        arm_num: 1
        edge: [1,2]
        type: release
        description: Put the apple into the drawer.
    node_4:
        name: task complete
        arm_num: 0
        edge: [3]
        type: complete
        description: Task complete after node_3 has been executed.
Now, let’s begin:
E.2 Baseline: Prompt for TP-S
{mdframed}
[backgroundcolor=azure, linewidth=0pt] You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot. To be on accuracy side, you can only use single arm or manipulate same object use dual arm each stage.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages. The (side) should never be omitted.
2) The stage name must adhere the format: API (side). The API must select in command API and follow the description.
3) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
4) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
{mdframed}[backgroundcolor=azure, linewidth=0pt] 5) The next stage you generate should accord to hand. For example, right_hand: holding "apple", left_hand: empty. The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand. The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
6) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
7) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
8) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages. For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and stage_next: grasp pot_right_handle (right).
9) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
10) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any stage that is not relevant to the completion of the task should not be present.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
  - switch_off <target-object>
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour <object> into <target_object>
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
You should only respond in the format as described below and do not respond anything else:
RESPONSE FORMAT:
stages:
[…]
        name: The name of the stage.
        right_hand_state: Empty or holding sth.
        left_hand_state: Empty or holding sth.
{mdframed}[backgroundcolor=azure, linewidth=0pt] Here’s an example input and response:
INPUT:
task: put "apple" into "drawer_floor"
obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints:  ["drawer_joint"]
      spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
        "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
        "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
    "table":
        type: actor
        fixed: True
        spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
        description: None
RESPONSE:
stages:
    stage_1:
        name: open "drawer_floor" (left)
        right_hand: empty
        left_hand: empty
        description: Open the drawer to access inside. This stage use left arm.
    stage_2:
        name: grasp "apple" (right)
        right_hand: holding "apple"
        left_hand: Grasp the apple on the table. This stage use right arm.
    stage_3:
        name: put "apple" into "drawer_floor" (right)
        right_hand: empty
        left_hand: empty
        description: Put the apple into the drawer use right arm. This stage use right arm.
Now, let’s begin:
E.3 Baseline: Prompt for TP-D
{mdframed}
[backgroundcolor=azure, linewidth=0pt]
You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot. To be on efficiency side, you can use dual arm each stage.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages.
2) The stage name should adhere the format: "API (right) and API (left)". The API must select in command API and follow the description. Note that right API and left API is executing at the same time, without order.
3) If you need to use both hand too manipulate same object, the stage name should adhere the format: "API (both)".
4) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
5) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
6) The next stage you generate should accroding to hand. For example, right_hand: holding "apple", left_hand: empty. The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand. The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
7) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
8) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
9) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages. For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and grasp pot_right_handle (right).
10) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
11) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any stage that is not relevant to the completion of the task should not be present.
12) Please try to take advantage of the dual-arm robot’s two-armed feature.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
  - switch_off <target-object>
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
{mdframed}[backgroundcolor=azure, linewidth=0pt]
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour
<object> into <target_object>
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
* The following command allows robot to do nothing:
  command syntax:
  - None
You should only respond in the format as described below and do not respond anything else:
RESPONSE FORMAT:
stages:
[…]
        name: The name of the stage.
        right_hand_state: Empty or holding sth.
        left_hand_state: Empty or holding sth.
Here’s an example input and response:
INPUT:
task: put "apple" into "drawer_floor"
obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints:  ["drawer_joint"]
     spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
         "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
    "table":
        type: actor
        fixed: True
{mdframed}[backgroundcolor=azure, linewidth=0pt]         spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
        description: None
RESPONSE:
stages:
    stage_1:
        name: grasp "apple" (right) and open "drawer_floor" (left)
        right_hand: holding "apple"
        left_hand: empty
       description: Grasp the apple on the table use right arm and open the drawer to access inside use left arm.
    stage_2:
        name: put "apple" into "drawer_floor" (right) and None (left)
        right_hand: empty
        left_hand: empty
        description: Put the apple into the drawer use right arm.
E.4 Insturction and Environment Description of Task5
{mdframed}
[backgroundcolor=azure, linewidth=0pt] task: heat the soup and put the tin on the table
constraints:
    1) do not need to take bowl out of microwave after heating
    2) close the door of refrigerator cooler after using
obj_names:["table", "dishwasher","dishwasher_rack", "dishwasher_door", "microwave","microwave_body", "microwave_door", "tin", "bowl", "refrigerator_cooler_body", "refrigerator_cooler_door"]
obj_description:
    "table":
        type: actor
        fixed: True
        spatial_relationship: The "table" is on the ground.
        description: None
    "dishwasher":
        type: articulation
        fixed: True
        links: ["dishwasher_shell", "dishwasher_rack", "dishwasher_door", "dishwasher_handle", "dishwasher_switch"
        joints: ["door_joint", "switch_joint"]
       spatial_relationship: The "dishwasher" is under the "table". The "dishwasher" is at the left part of the "table".
        description:
            "dishwasher_shell": The shell of the "dishwasher".
            "dishwasher_rack": The rack of the "dishwasher".
            "dishwasher_door": The door of the "dishwasher".
            "microwave_handle": The handle of the "dishwasher".
            "dishwasher_switch": The switch of the "dishwasher".
            "door_joint": The "door_joint" enables opening and closing of "dishwasher_door" by manipulating "microwave_handle".
            "switch_joint": The "switch_joint" enables and disables of "dishwasher" by switching off and switching on "dishwasher_switch". The "dishwasher_switch" can only be manipulated when "dishwasher_door" is closing.
    "microwave":
        type: articulation
        fixed: True
[…]

### v3

@@ IV-B Evaluation of Planning Effectiveness and Conciseness
TP-S $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}9.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}53.5}=63.0$ $40\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}12.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}90.7}=103.6$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.6}=43.9$ $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}116.9}=125.1$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}55.8}=64.0$ $78\%$ $100.0\%$
TP-D $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.1}=34.6$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=51.4$ $26\%$ $111.4\%$
TwoStep[11] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.3}=38.7$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=54.5$ $26\%$ $109.2\%$
RoCo[12] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}49.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.2}=75.5$ $0\%$ $Fail$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.1}=79.0$ $30\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}155.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}78.9}=234.2$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.3}=89.7$ $60\%$ $68.7\%$
DAG-Plan $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.6}=36.1$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.7}+{\color[rgb]{0.78515625,0.11328125,0.19140625}59.8}=71.5$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.0}=32.3$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}69.9}=77.9$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.4}=43.7$ $74\%$ $152.8\%$
As shown in Table II, in the plan tests, DAG-Plan consistently outperformed baselines, showcasing superior efficiency and robustness. DAG-Plan achieved a high success rate across all tasks, demonstrating its effectiveness in dual-arm manipulation. Notably, it maintained an impressive macro average success rate of 88% and exhibited a significant reduction in the required stages for task completion, achieving 169.2% stage efficiency. These results underscore the effectiveness of DAG-Plan in achieving task goals and its efficiency in execution.
In contrast, TP-S, primarily focused on the single-arm plan, generally required more stages to complete tasks compared to TP-D and DAG-Plan. Although TP-S maintained a relatively high and consistent success rate, it was less efficient in stage minimization.
Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 40%), and often produced plans that were not executable in the physical environment. While theoretically capable of reducing the number of stages required for tasks, TP-D frequently encountered challenges related to coordination complexities and unrealistic task assignments. This highlights the superiority of DAG-Plan in effectively translating high-level plans into physical actions and navigating complexities in the environment with relative ease compared to TP-D.
The key idea of Twostep is to decompose a task into two completely independent sub-tasks. While this approach is highly effective in multi-agents task settings, it falls short in dual-arm tasks. In dual-arm scenarios, the entire task sequence is interconnected, especially in tasks with order requirements, such as Task 3 and Task 4. Each sub-task is inherently dependent on others, making it impossible to decompose the task into fully independent sub-tasks. As a result, Twostep achieves low success rates and is ineffective in dual-arm tasks.
RoCo is capable of acquiring the current environment and robot state at each step. Through communication between the left and right arms, it can often select appropriate subtasks, achieving high success rates and phase efficiency for all tasks except Task 2. In Task 2, RoCo is still limited by the drawback of sequentially generating dual-arm plans, neglecting implicit dependency conditions. For instance, it attempts to grasp the sponge and mug without first opening the drawer, resulting in a failure to open the drawer and ultimately causing the plan to fail. Additionally, RoCo requires querying the large model and retaining historical dialogue in each round, leading to significant token consumption in long-sequence tasks.
The success of DAG-Plan can be attributed to its DAG generation, which systematically breaks down tasks into manageable components. This allows for more accurate and feasible plan generation, as the method iteratively refines the task execution steps while maintaining alignment with the operational capabilities of the robotic system. The failed case of DAG-Plan is LLMs occasionally generate some unreasonable task graphs. For instance, each node depend on the previous node, causing the task graph to degenerate into a task sequence, or certain dependencies be omitted.
IV-C Evaluation with Physical Simulation
[…]